Before the year​ began, McCrae, Inc. estimated its manufacturing overhead costs for the year to be ​$300,000 and estimated that 20,000 machine hours would be logged during production. The company uses machine hours to apply manufacturing overhead to production. During the​ year, the company received an order from a customer for 15 identical units of its most popular product.​ $1,000 in direct materials were requisitioned and​ $1,500 in direct labor costs were incurred during the production of the units. A total of 14 machine hours were logged during the production of the units. Assuming a sales price of ​$558​, the gross profit reported on each unit in the order equals ​$____.​
Answer #1

manufacturing overhead per machine hour = 300000/20000 = 15 per machine hour

cost of production = direct material + direct labor + manufacturing overhead

= 1000+1500+(14*20) = 2780

gross profit = sales- cost of goods produced

(15*558)-2780 = 5590

gross profit per unit = 5590/15 = 372.67